OpenStudy (anonymous):

From a hot-air balloon directly over a target, you throw a marker with an initial velocity of -40 feet per second from a height of 180 feet. How long does it take the marker to reach the target?

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (anonymous):

Can anybody help me!

OpenStudy (aum):

s = ut + 1/2 * g * t^2 s = 180 ft u = 40 feet / sec g = 32 ft / sec^2 Solve for t.

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

@aum

OpenStudy (aum):

Yes? Just plug the numbers into the equation and solve for t.

OpenStudy (aum):

s = ut + 1/2 * g * t^2 s = 180 ft; u = 40 feet / sec; g = 32 ft / sec^2 180 = 40t + 1/2 * 32 * t^2 180 = 40t + 16t^2 16t^2 + 40t - 180 = 0 4t^2 + 10t - 45 = 0 This is a quadratic equation that you can solve using the quadratic formula.

OpenStudy (anonymous):

so like this 180=40t+1 2 ⋅32 2 ⋅t 2

OpenStudy (aum):

4t^2 + 10t - 45 = 0 Compare it to the general quadratic formula: ax^2 + bx + c = 0 a = 4, b = 10, c = -45 Plug it into the quadratic formula: \(\Large \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a } \)

OpenStudy (anonymous):

is the answer 90ft./sec

OpenStudy (anonymous):

@aum

OpenStudy (aum):

I don't think so. But the answer will be in seconds (not ft. / sec) as the question asks for "how long" will it take which refers to time. You can plug in the numbers into the quadratic equation and solve for t or you can graph using either a graphing calculator or an online tool and find where it crosses the x-axis.

OpenStudy (anonymous):

venny can u give me an answer 2 tha ques

OpenStudy (anonymous):

@magbak hey bae I missed u

OpenStudy (anonymous):

did u read my testimony I sent u

OpenStudy (anonymous):

ok? I guess no response then, huh

OpenStudy (anonymous):

@magbak